# PEGASUS: Physically Enhanced Gaussian Splatting Simulation System for 6DOF Object Pose Dataset Generation

### v1

@@ IV-A3 Physics Engine
We selected PyBullet [39] as our physics engine, renowned for its lightweight and versatile capabilities. Within the simulation, the environment mesh is integrated as a static component. For the objects, we determine an appropriate height to drop varying quantities (user-selected) into the scene, simulating natural object placement. Throughout the simulation, we track and extract the orientation and translation of each object, represented as a quaternion and a translation vector, respectively. PEGASUS is thus equipped to simulate both static and dynamic scenes, offering a comprehensive range of possibilities for scene creation and analysis.
@@ IV-B PEGASUS Dataset Generation
To create a dataset using Gaussian Splatting with PEGASUS, we integrate the Gaussian Splatting base environment and Gaussian Splatting objects. This approach, enhanced by physical placement techniques, allows us to generate an unlimited number of unique scenes. To generate a new scene, we simulate the trajectory using our physics engine, apply transformations to each Gaussian Splatting object, and render the scene with the Gaussian Splatting rasterizer [12].
In the context of 6DoF object pose estimation using this dataset generation tool, we extract the transformations of objects from the model coordinate system to the camera coordinate system. This process determines the pose of the objects within the scene. To render the scene from various viewpoints, we select a set of random ground truth poses and create a trajectory by interpolating between these poses.
Ultimately, we save the raw data, including rendered RGB images, depth maps, segmentation maps of the silhouette and visible masks, 2D/3D bounding boxes as well as the transformation matrices from object to world and world to camera coordinate system. This data is formatted according to the popular BOP-dataset format [30].
We want to emphasise that PEGASUS is also capable of rendering dynamic scenes, which broadens the potential applications of the generated data. It is worth noting that this approach can be easily extended with custom data. To incorporate a new environment, one only needs to record a set of images and convert them into a Gaussian Splatting instance. The same process applies to any object; a simple scan is enough to compute the photometric and geometric properties required for integration into PEGASUS. A set of images extracted from the dataset generator is shown in Fig.3.
@@ IV-C Data Set
In this section, we discuss the Ramen dataset and outline the pipeline for processing each individual object. In an earlier work [28] a similar scanning system was used to create NeRF representations of household objects.
The Ramen dataset comprises over 30 varieties of cup noodles, readily available in most mini markets. A comprehensive summary of all products is illustrated in Fig. 4.
For recording the objects, we employed the 3D PhotoBench 280 from Ortery [40], a commercial $360^{\circ}$ turntable system, alongside the 3D MultiArm 2000 [41] camera system. This setup enabled synchronous image acquisition from five different angles capturing 150 images for each hemisphere and automatic background removal.
The initial step involved scanning a planar calibration board with a feature-rich surface and an ArUco marker of known size. This process, aided by COLMAP [19], allowed for precise pose computation. Subsequently, the scene (including poses and point cloud) was scaled to achieve a metric reconstruction [33]. Scene alignment was conducted by fitting a plane into the point cloud using RANSAC [36], ensuring the normal vector of the visible plane faced the positive $\mathbf{z}$-direction. The poses obtained from the calibration board were repurposed as a calibration reconstruction.
For each cup noodle product, the poses from the upper hemisphere’s calibration reconstruction were utilized. The sparse point cloud was then recomputed by triangulating the matched feature points from the product images. To capture photometric information about the bottom part, we scanned the bottom hemisphere of the flipped product in a similar fashion. However, it was not possible to reuse the calibration target for the bottom hemisphere, as the flipping of the object altered the pose locations. Our methodology involved registering the bottom images into the existing top reconstruction, resulting in approximately 270 registered images per product. The bottom row remained unregistered due to its texture-less appearance, however this did not seem to affect the quality of the reconstruction.
Each object was reconstructed using 3DGS and a low-poly mesh extracted from the point cloud (as detailed in section IV-A2). To integrate these meshes into the physics engine, we generated a URDF file for every object (including environment objects), documenting both visual and collision model information.
@@ V Experiments
This section presents experiments demonstrating the successful application of data generated by PEGASUS in training a neural network. Additionally, it showcases the use of a UR5 robot to execute real-world pick-and-place operations.
We selected the Deep Object Pose (DOPE) [42] network structure for our experiments. For dataset generation, three distinct data sets were created, each comprising 60,000 images featuring a single cup noodle, set in three different environments. In total, we generated 2,000 unique scenes with 30 images per scene, captured from various perspectives. The generation process took 6 hours on a laptop with an Intel i9 12th Gen CPU and NVIDIA RTX 3080 Ti (Mobile) GPU.
Regarding training, we utilized the default hyper-parameters of DOPE and trained the network for 15 epochs. The UR5 robot was configured to accurately pick the center of the cup noodle and place it into a basket. Our experiments successfully demonstrated the capability of the robot to sequentially pick up 10 out of 10 cup noodles in a row, as well as to grasp various types of cup noodles.
@@ VI Limitations
Our method, while effective, is not without its limitations. One significant shortfall is the absence of realistic shadow rendering in our system. Consequently, incorporating shadow maps or screen space ambient occlusion represents a natural and necessary next step in our development process. Additionally, when placing objects within an environment, our current approach does not account for re-lighting, scattering, refraction, or reflection. This omission can result in scenes that appear somewhat unnatural.
Another challenge we faced involves scanning texture-less environments, which often leads to a noisy Gaussian splatting reconstruction. This noise manifests as large Gaussian splats that may overlap or interfere with objects, potentially causing visual artifacts. Addressing these issues is crucial for enhancing the realism and visual fidelity of our rendered scenes.
@@ VII Conclusion
We have introduced PEGASUS, a versatile dataset generator designed to enhance accuracy and quality in object pose estimation. Alongside PEGASUS, we present the Ramen dataset, which includes over 30 diverse products. The dataset generator adeptly creates photorealistic renderings, semantic masks, depth maps, and captures the object pose. PEGASUS is specifically engineered to generate domain-specific data sets, aiding in the fine-tuning of neural networks that extend beyond mere pose estimation tasks.
To further empower PEGASUS, it is crucial to accumulate a more extensive collection of environments and objects. This expansion is key to evolving towards a more generalizable dataset generator. Another intriguing avenue for enhancement involves applying augmentations directly to the objects or their environments. Techniques like diffusion models, such as GaussianDreamer [18] or Rosie [17], can be employed to alter the shape and appearance of objects and environments, offering a new dimension of flexibility in dataset generation.
@@
Acknowledgement
We extend our sincere gratitude to Abdullah Mustafa for his valuable feedback and to Toshio Ueshiba for his extensive expertise in UR5. This paper is one of the achievements of joint research with and is jointly owned copyrighted material of ROBOT Industrial Basic Technology Collaborative Innovation Partnership. This research has been supported by the New Energy and Industrial Technology Development Organization (NEDO), under the project ID JPNP20016.

### v2

@@ IV-A3 Physics Engine
We selected PyBullet [39] as our lightweight physics engine. Within the simulation, the environment mesh is integrated as a static component. For the objects, we determine an appropriate height to drop varying quantities (user-selected) into the scene, simulating natural object placement. Throughout the simulation, we track and extract the orientation and translation of each object, represented as a quaternion and a translation vector, respectively. PEGASUS is thus equipped to simulate both static and dynamic scenes, offering a comprehensive range of possibilities for scene creation and analysis.
@@ IV-B PEGASUS Dataset Generation
To create a dataset using Gaussian Splatting with PEGASUS, we include Gaussian Splatting base environments and Gaussian Splatting objects. This approach, enhanced by physical placement techniques, allows us to generate an unlimited number of unique scenes. To generate a new scene, we simulate the trajectory using our physics engine, apply transformations to each Gaussian Splatting object, and render the scene with the Gaussian Splatting rasterizer [9].
To render the scene from various viewpoints, we randomly select a set of ground truth poses and create a trajectory by interpolating between these poses. Lastly, we save the raw data, including rendered RGB images, depth maps, segmentation maps of the silhouette and visibility masks, 2D/3D bounding boxes as well as the transformation matrices from object to world and world to camera coordinate systems. This data is formatted according to the popular BOP-dataset format [10]. We want to emphasize that PEGASUS is also capable of rendering dynamic scenes, which broadens the potential applications of the generated data. It is worth noting that this approach can be easily extended with custom data. To incorporate a new environment, one only needs to record a set of images and convert them into a 3DGS instance. The same process applies to any object; a simple scan is enough to compute the photometric and geometric properties required for integration into PEGASUS. A set of images extracted from the dataset generator is shown in Fig. 3.
@@ IV-C Data Set
The focus of our research is on the development of robotic systems in the service sector to support personnel in retail. With Japan having one of the highest population densities, it boasts a considerable number of 24-hour convenience stores known as konbini [40]. Introducing robots into these compact retail spaces can prove to be a valuable asset for tasks like restocking products and efficiently managing inventory. Therefore we focused on the specific product group of ramen noodles gathered in the Ramen dataset. The dataset comprises of 30 varieties of cup noodles, readily available in most mini markets. A comprehensive summary of all products is illustrated in Fig. 4.
For recording the objects, we employed the 3D PhotoBench 280 from Ortery [41], a commercial $360^{\circ}$ turntable system, alongside the 3D MultiArm 2000 [42] camera system. This setup enabled synchronous image acquisition from five different angles capturing 150 images for each hemisphere and applying automatic background removal. The process is heavily inspired by Neural Scanning [26].
The initial step involved scanning a planar calibration board with a feature-rich surface and an ArUco marker of known size. This process, aided by COLMAP [18], allowed for precise pose computation. Subsequently, the scene (including poses and point cloud) was scaled to achieve a metric reconstruction [30] and aligned to the $xy$ plane ensuring the normal vector of the visible plane faced the positive $\mathbf{z}$-direction. The poses obtained from the board were repurposed as a calibration reconstruction. For each cup noodle product, the poses from the upper hemisphere’s calibration reconstruction were utilized. The sparse point cloud was then recomputed by triangulating the matched feature points from the product images. To capture photometric information about the bottom part, we scanned the bottom hemisphere of the flipped product symmetrically. However, it was not possible to reuse the calibration target for the bottom hemisphere, as the flipping of the object altered the pose locations. Our methodology involved registering the bottom images into the existing top reconstruction, resulting in approximately 270 registered images per product. Each object was reconstructed using 3DGS and a low-poly mesh was extracted from the point cloud (as detailed in section IV-A2). To integrate these meshes into the physics engine, we store the information for every object, including environment objects, in a Unified Robotics Description Format (URDF) file, documenting both visual and collision model information.
In addition to the Ramen dataset, we provide PEGASET as a second dataset. It comprises a selection of 21 objects from the well-known YCB-V Dataset [1]. By including these fundamental objects, we aim to encourage broader adoption by the robotics community.
@@ V Experiments
This section presents experiments demonstrating the successful use of Universal Robots UR5 to execute real-world pick-and-place operations on data generated by PEGASUS.
We selected the Deep Object Pose (DOPE) [11] network structure for our experiments. For dataset generation, three distinct data sets were created, each comprising 60,000 images featuring a single cup noodle, set in three different environments. In total, we generated 2,000 unique scenes with 30 images per scene, captured from various perspectives. The generation process took 6 hours on a laptop with an Intel i9 12th Gen CPU and NVIDIA RTX 3080 Ti (Mobile) GPU. Regarding training, we utilized the default hyper-parameters of DOPE and trained the network for 15 epochs. The UR5 was configured to accurately pick the center of the cup noodle and place it into a basket. Our experiments successfully demonstrated the capability of the robot to sequentially pick up 10 out of 10 cup noodles in a row, as well as to grasp various types of cup noodles.
@@ VI Limitations
Our method, while effective, is not without its limitations. One significant shortfall is the absence of realistic shadow rendering in our system. Consequently, incorporating shadow maps or screen space ambient occlusion represents a natural and necessary next step in our development process. Additionally, when placing objects within an environment, our current approach does not account for re-lighting, scattering, refraction, or reflection. This omission can result in scenes that appear somewhat unnatural. Another challenge we faced involves scanning texture-less environments, which often leads to a noisy Gaussian splatting reconstruction. This noise manifests as large Gaussian splats that may overlap or interfere with objects, potentially causing visual artifacts. Addressing these issues is crucial for enhancing the realism and visual fidelity of our rendered scenes.
@@ VII Conclusion
We have introduced PEGASUS, a versatile dataset generator designed to enhance accuracy and quality in object pose estimation. Alongside PEGASUS, we present the Ramen dataset, which includes over 30 diverse products. The dataset generator adeptly creates photorealistic renderings, semantic masks, and depth maps, and captures the object pose. PEGASUS is specifically engineered to generate domain-specific data sets, aiding in the fine-tuning of neural networks that extend beyond mere pose estimation tasks.
To further empower PEGASUS, it is crucial to accumulate a more extensive collection of environments and objects. This expansion is key to evolving toward a more generalizable dataset generator. Another intriguing avenue for enhancement involves applying augmentations directly to the objects or their environments. Techniques like diffusion models, such as GaussianDreamer [17] or Rosie [16], can be employed to alter the shape and appearance of objects and environments, offering a new dimension of flexibility in dataset generation.
@@
Acknowledgment
We extend our sincere gratitude to Abdullah Mustafa for his valuable feedback and to Toshio Ueshiba for his extensive expertise in UR5.